The Moon has a diameter of approximately 3.476×103 m. Assuming the Moon is a 

sphere, calculate the surface area in square kilometers


1
Expert's answer
2021-07-25T09:18:26-0400
"A=4\pi r^2=4\pi\cdot(3.476\cdot10^3\ m)^2=1.52\cdot10^8\ m^2,""A=1.52\cdot10^8\ m^2\cdot\dfrac{1\ km^2}{10^6\ m^2}=152\ km^2."